OpenStudy (anonymous):

What is the slope, and what is the y-intercept of the line? y=1/2x-5 slope = -------------------------------------------------------------------------------- y-intercept =

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

y = mx+b M=slope B=y int

OpenStudy (dmndlife24):

First you would need to rearrange the equation in y = mx + b format (better known as slope - intercept form). But since this equation is already in slope-int. form, you do not have to worry about this step. Our equation is y = 1/2x - 5 m = slope and b = y - int. m = slope = 1/2 b = y - intercept = -5 (0, -5)